# Reward Engineering for Spatial Epidemic Simulations: A Reinforcement Learning Platform for Individual Behavioral Learning

### v1

@@ Abstract
We present ContagionRL, a Gymnasium-compatible reinforcement learning platform specifically designed for systematic reward engineering in spatial epidemic simulations. Unlike traditional agent-based models that rely on fixed behavioral rules, our platform enables rigorous evaluation of how reward function design affects learned survival strategies across diverse epidemic scenarios. ContagionRL integrates a spatial SIRS+D epidemiological model with configurable environmental parameters, allowing researchers to stress-test reward functions under varying conditions including limited observability, different movement patterns, and heterogeneous population dynamics. We evaluate five distinct reward designs, ranging from sparse survival bonuses to a novel potential field approach, across multiple RL algorithms (PPO, SAC, A2C). Through systematic ablation studies, we identify that directional guidance and explicit adherence incentives are critical components for robust policy learning. Our comprehensive evaluation across varying infection rates, grid sizes, visibility constraints, and movement patterns reveals that reward function choice dramatically impacts agent behavior and survival outcomes. Agents trained with our potential field reward consistently achieve superior performance, learning maximal adherence to non-pharmaceutical interventions while developing sophisticated spatial avoidance strategies. The platform’s modular design enables systematic exploration of reward-behavior relationships, addressing a knowledge gap in models of this type where reward engineering has received limited attention. ContagionRL is an effective platform for studying adaptive behavioral responses in epidemic contexts and highlight the importance of reward design, information structure, and environmental predictability in learning.
@@ 1 Introduction
The dynamics of infectious outbreaks, epidemics or pandemics, are intricately linked to individual behavior (wang_2021; funk_2010). Behavioral decisions such as movements, risk perceptions and adherence to non-pharmaceutical interventions (such as mask-wearing and social distancing) could heavily affect how an epidemic unfolds (wang_2021; funk_2010; manrubia_2022; ye_2020; saad_npi_2023). Traditional epidemic models primarily use ordinary differential equations (ODEs) to keep track of disease population-wide with models such as SIR (kermack1927contribution) or SEIR (kermack1927contribution; wilson1945law). While these models are effective for a high-level overview of the disease outbreak, due to their assumption of homogeneity, they neglect nuanced effects of individual behavior and spatial interactions (bansal_2007; zachreson_2022). Furthermore, they are incapable of capturing dynamic behavioral changes or adaptive interventions (funk_2010; manrubia_2022; bansal_2007).
@@
The SARS-CoV-2 pandemic highlighted how individual-level decisions and behaviors, such as adherence to non-pharmaceutical interventions (NPIs) like mask wearing, could profoundly change the trajectory of a disease outbreak (saad_npi_2023). Agent-Based Models (ABMs) (grimm_2013_abm), also known as Individual-Based Models (IBMs), offer a compelling alternative by simulating systems from the bottom up, representing individual agents with specific attributes, spatial contexts, and behavioral rules. These models accommodate heterogeneity and complex interactions, thereby making ABMs particularly suitable for studying epidemics where individual actions drive epidemic-scale outcomes, while leaving them reliant on prescribed or rule-based behaviors rather than learned policies that exhibit more complexity and are more dynamic (grimm_2013_abm).
@@
The lack of the learned policies could be addressed with Reinforcement Learning (RL) (Sutton_Barto_textbook) and Deep Reinforcement Learning (DRL) (li_2017_DRL; mousavi_2016_DRL; franccois_2018_DRL) that provide powerful frameworks to optimize sequential decision making in dynamic systems (mnih_2015; silver_2016; arulkumaran_2017) by modeling them as Markovian decision processes (MDP). Such algorithms have been applied in epidemiological contexts to learn the optimal policies by focusing on centralized control of macroscopic interventions (libin_2021; kwak_2021; HRL4EC_du_2023; Wan_2021). Combining of RL or DRL algorithms with ABMs allows simulation of both adaptive and heterogeneous behaviors.
@@
However, the main challenge in creating and developing informative ABMs lies in defining realistic and computationally tractable agent behaviors. Furthermore, RL-based approaches are highly sensitive to the reward function and the reward design can have considerable impact on the learned policy and the downstream interpretations of the model upon behavioral convergence (ng1999policy; importance_of_reward_design_in_RL_2024; importance_of_reward_design_in_RL_2025; shihab2025detectingmitigatingrewardhacking).
@@ 2 Related Works
Compartmental Models in Epidemiology. Classical epidemiological models, such as SIR (Susceptible-Infected-Recovered), divide a population into discrete groups (compartments) based on disease state (kermack1927contribution; wilson1945law). The flow of individuals between compartments is described by ordinary differential equations (ODEs). Extensions like SEIR (adding an exposed group to act as a disease incubation period) or SVE(R)IRS (adding Vaccinated, accounting for waning immunity) incorporate more states to increase realism (brauer2019mathematical; friedman2014mathematical; hethcote2000mathematics; Chyba_2024; Balisacan_2021). While traditional compartmental models provide valuable population-level insights, assumptions of homogeneity limit their ability to represent individual decision-making (Cangiotti_2024; Siegenfeld_2022; Dimarco_Perthame_Toscani_Zanella_2021; Jiang2024ModelingIA; Bostanci_Conrad_2025; Rose_Medford_Goldsmith_Vegge_Weitz_Peterson_2021; Kong_Wang_Han_Cao_2016). These models also struggle to capture dynamic behavioral changes or adaptive interventions (brauer2019mathematical; Siegenfeld_2022). Ignoring such factors can limit accuracy and lead to misleading conclusions (Sudhakar_2024; Siegenfeld_2022). Stochastic formulations are particularly relevant for small populations or during the early stages of an outbreak (brauer2019mathematical).
@@
Agent-Based Models (ABMs). ABMs address the limitations of compartmental models like SIR by modeling heterogeneous individuals using specific attributes and interaction rules within an environment or network (Tracy_Cerdá_Keyes_2018; Hunter_Mac_Namee_Kelleher_2018; Hunter_2017). Furthermore, ABMs excel at capturing individual variability, stochasticity, local interactions, and complex contact networks, which are difficult for compartmental models (Siegenfeld_2022; Hunter_2017). Recent advances demonstrate substantial sophistication such as large-scale ABMs that can simulate millions of agents with GPU acceleration (chopra2021deepabm) or employ LLM-based adaptive agents (park2023generative).
@@
Challenge of Reward Engineering. Sophisticated ABMs show potential for learning-enabled epidemic modeling. However, systematic evaluation of reward function design remains an important methodological gap. In reinforcement learning, reward design shapes the fundamentals of the learned behavior with different reward structures leading to considerably different learned policies at convergence (ng1999policy; amodei2016concrete; importance_of_reward_design_in_RL_2025; shihab2025detectingmitigatingrewardhacking). Despite the importance of this problem in RL, epidemic modeling applications have not systematically investigated how reward structure affects learning under varying conditions.
@@
Reinforcement Learning (RL) for Epidemics. RL has been increasingly leveraged to design policies for epidemic intervention at population and individual level. Individual-based Reinforcement Learning Epidemic Control Agent (IDRLECA) proposed an RL agent at an individual level using GNNs to estimate infection probabilities and balance epidemic suppression vs cost of mobility in its reward function (feng2023contact). VEHICLE (feng2022precise) handles challenges of unobservable asymptomatic cases and delayed effects by using a combination of hierarchical RL with GNNs. Recent work has also focused on using actor critic methods to minimize a cost that considers both the epidemiological, economic and social cost for multi-intervention planning (Mai_2023) or using RL for scheduling lockdown periods (arango2020covid) that incorporates the ICU capacity and minimizing time spent in lockdowns. HRL4EC (HRL4EC_du_2023) also leverages hierarchical RL and multi-component reward that considers both health and socioeconomic cost. SiRL (Bushaj_2023) uses a detailed ABM with RL and feeds the compartmental levels to the agent and uses rewards that are tied to intervention effectiveness and disease spread. These approaches excel at macro-level policy optimization such as determining lockdown time, resource allocation, and intervention sequencing (Mai_2023; arango2020covid; khadilkar2020optimising). While valuable for analyzing epidemic policy, these centralized approaches do not examine the impact of different reward formulations on the learned RL policy.
@@ 3 Methodology
ContagionRL is a custom Gymnasium-compatible (Gymanasium) environment that integrates a SIRS+D compartmental epidemiological model (hethcote2000mathematics). Built within an agent-based modeling framework, the environment simulates a single reinforcement learning (Sutton_Barto_textbook) agent interacting with a population of humans whose behaviors may be stochastic or deterministic or mix of both. This design enables the study of learned risk-averse navigation and adherence behaviors in response to an evolving epidemic (saad_npi_2023) in different contexts.
Single Agent Design Rationale. Epidemic dynamics are inherently multi-agent interactions (manfredi2013modeling; wang2016statistical). However, our single agent approach is a methodological choice that enables us to evaluate the design of the reward function in a systematic format. Multi-agent environments introduce complexities such as non-stationarity due to other learning agents (foerster2017stabilising), emergent social behavior (leibo2017multi) and strategic interactions (tampuu2017multiagent). Each of these complexities can complicate and potentially obfuscate how specific reward components influence the learned policy at an individual level. The isolation of the learning process to a single agent in a controlled population of non-learning humans, allows for transparent study of the impact of reward structure. This permits the identification of essential reward components for effective behavior learning. Furthermore, this presents a first step for understanding the relationship between reward and behavior before applying insights to a complex domain like multi-agent reinforcement learning. By isolating a single learning agent, in a controlled population of non-learning humans, we can conduct detailed assessments of how different reward formulations affect the learned strategies without interference from adaptive social behaviors. The controlled nature of this methodology allows insights into the relationship between reward and behavior, that will be transferable to more complex multi-agent deployments where understanding individual level incentive responses becomes critical for designing effective reward systems across a population.
@@ 3.1 SIRS+D Simulation Environment
Environment Dynamics. The simulation happens over discrete timesteps within a two-dimensional grid space of size $G\times G$ with periodic boundary conditions (periodic_boundary_computersim; periodic_boundary_animals; periodic_boundary_molecular). The environment is populated by $N$ non-learning humans and a single controllable Reinforcement Learning (RL) agent. Each individual, human or agent, can exist in one of four states: Susceptible (S), Infected (I), Recovered (R), or Dead (D). We conduct a set of sensitivity analysis in Appendix F and vary the infection rate of disease ($\beta$) in Figure 12 and Table 11, vary the grid size and population density in Figure 13 and Table 12, vary the adherence effectiveness ($\alpha_{\text{eff}}$) in Figure 14 and Table 13, and vary the distance decay of disease transmission in Figure 15 and Table 14.
@@ 3.2 RL Agent Formulation
We implement the agent’s learning problem as a configurable decision-making process that supports both Markov Decision Process (MDP) and Partially Observable MDP (POMDP) formulations. We use both the MDP and POMDP formulation in our experiments. The POMDP employs a threshold-based visibility constraint, where agents can only observe other entities within a specified Euclidean distance radius. The agent’s policy $\pi(a_{t}|o_{t})$, mapping observations to actions, is learned using three deep reinforcement learning algorithms: Proximal Policy Optimization (PPO) (schulman2017proximal), Soft Actor-Critic (SAC) (sac_algorithm), and Advantage Actor Critic (A2C) a synchronous variant of A3C algorithm (a3c_paper_mnih), all implemented via the stable-baselines3 library (stable-baselines3).
@@
Action Space. At each timestep $t$, the agent selects a continuous-valued action $a_{t}=(\Delta x,\Delta y,\alpha)$, where $(\Delta x,\Delta y)\in[-1,1]^{2}$ defines a movement vector and $\alpha\in[0,1]$ specifies the adherence level to non-pharmaceutical interventions (NPIs). NPIs are methods used to reduce the spread of an epidemic disease without any pharmaceutical drug treatments such as wearing masks or social distancing and are considered extremely effective ways of controlling primary outbreaks that reduce disease transmission and consequentially mortality (npi_effectiveness_2021; ferguson2020report). The agent’s position is updated using the movement vector, scaled by a maximum step size, and wrapped around the grid’s periodic boundaries:
@@ 4 Baselines
Greedy Distance Maximizer Baseline. This heuristic policy implements a specific survival strategy focused on immediate risk avoidance. At each step, it first identifies the nearest infected human. It then evaluates a predefined set of potential movement actions across the eight cardinal and intercardinal directions, plus staying stationary. It systematically simulates the outcome of a each potential action, by computing the hypothetical distance to the nearest infected human if that move were taken, and selects the action that yields the greatest distance separation. The policy’s evaluation of counterfactual outcomes provides a situational awareness absent in simpler baselines. While myopic and handcrafted, this policy provides a strong heuristic benchmark by explicitly evaluating the spatial consequence of each potential action.

### v2

@@ Abstract
We present ContagionRL, a Gymnasium-compatible reinforcement learning platform specifically designed for systematic reward engineering in spatial epidemic simulations. Unlike traditional agent-based models that rely on fixed behavioral rules, our platform enables rigorous evaluation of how reward function design affects learned survival strategies across diverse epidemic scenarios. ContagionRL integrates a spatial SIRS+D epidemiological model with configurable environmental parameters, allowing researchers to stress-test reward functions under varying conditions including limited observability, different movement patterns, and heterogeneous population dynamics. We evaluate five distinct reward designs, ranging from sparse survival bonuses to a novel potential field approach, across multiple RL algorithms (PPO, SAC, A2C). Through systematic ablation studies, we identify that directional guidance and explicit adherence incentives are critical components for robust policy learning. Our comprehensive evaluation across varying infection rates, grid sizes, visibility constraints, and movement patterns reveals that reward function choice dramatically impacts agent behavior and survival outcomes. Agents trained with our potential field reward consistently achieve superior performance, learning maximal adherence to non-pharmaceutical interventions while developing sophisticated spatial avoidance strategies. The platform’s modular design enables systematic exploration of reward-behavior relationships, addressing a knowledge gap in models of this type where reward engineering has received limited attention. ContagionRL is an effective platform for studying adaptive behavioral responses in epidemic contexts and highlight the importance of reward design, information structure, and environmental predictability in learning. Our code is publicly available at https://github.com/redradman/ContagionRL.
@@ 1 Introduction
The dynamics of infectious outbreaks, epidemics or pandemics, are intricately linked to individual behavior (Wang et al., 2021; Funk et al., 2010). Behavioral decisions such as movements, risk perceptions and adherence to non-pharmaceutical interventions (such as mask-wearing and social distancing) could heavily affect how an epidemic unfolds (Wang et al., 2021; Funk et al., 2010; Manrubia and Zanette, 2022; Ye et al., 2020; Saad-Roy and Traulsen, 2023). Traditional epidemic models primarily use ordinary differential equations (ODEs) to keep track of disease population-wide with models such as SIR (Kermack and McKendrick, 1927) or SEIR (Kermack and McKendrick, 1927; Wilson and Worcester, 1945). While these models are effective for a high-level overview of the disease outbreak, due to their assumption of homogeneity, they neglect nuanced effects of individual behavior and spatial interactions (Bansal et al., 2007; Zachreson et al., 2022). Furthermore, they are incapable of capturing dynamic behavioral changes or adaptive interventions (Funk et al., 2010; Manrubia and Zanette, 2022; Bansal et al., 2007).
@@
The SARS-CoV-2 pandemic highlighted how individual-level decisions and behaviors, such as adherence to non-pharmaceutical interventions (NPIs) like mask wearing, could profoundly change the trajectory of a disease outbreak (Saad-Roy and Traulsen, 2023). Agent-Based Models (ABMs) (Grimm and Railsback, 2013), also known as Individual-Based Models (IBMs), offer a compelling alternative by simulating systems from the bottom up, representing individual agents with specific attributes, spatial contexts, and behavioral rules. These models accommodate heterogeneity and complex interactions, thereby making ABMs particularly suitable for studying epidemics where individual actions drive epidemic-scale outcomes, while leaving them reliant on prescribed or rule-based behaviors rather than learned policies that exhibit more complexity and are more dynamic (Grimm and Railsback, 2013).
@@
The lack of the learned policies could be addressed with Reinforcement Learning (RL) (Sutton and Barto, 2018) and Deep Reinforcement Learning (DRL) (Li, 2017; Mousavi et al., 2016; François-Lavet et al., 2018) that provide powerful frameworks to optimize sequential decision making in dynamic systems (Mnih et al., 2015; Silver et al., 2016; Arulkumaran et al., 2017) by modeling them as Markovian decision processes (MDP). Such algorithms have been applied in epidemiological contexts to learn the optimal policies by focusing on centralized control of macroscopic interventions (Libin et al., 2021; Kwak et al., 2021; Du et al., 2023; Wan et al., 2021). Combining of RL or DRL algorithms with ABMs allows simulation of both adaptive and heterogeneous behaviors.
@@
However, the main challenge in creating and developing informative ABMs lies in defining realistic and computationally tractable agent behaviors. Furthermore, RL-based approaches are highly sensitive to the reward function and the reward design can have considerable impact on the learned policy and the downstream interpretations of the model upon behavioral convergence (Ng et al., 1999; Ibrahim et al., 2024; Nguyen et al., 2025; Shihab et al., 2025).
@@ 2 Related Works
Compartmental Models in Epidemiology. Classical epidemiological models, such as SIR (Susceptible-Infected-Recovered), divide a population into discrete groups (compartments) based on disease state (Kermack and McKendrick, 1927; Wilson and Worcester, 1945). The flow of individuals between compartments is described by ordinary differential equations (ODEs). Extensions like SEIR (adding an exposed group to act as a disease incubation period) or SVE(R)IRS (adding Vaccinated, accounting for waning immunity) incorporate more states to increase realism (Brauer et al., 2019; Friedman and Kao, 2014; Hethcote, 2000; Chyba et al., 2024; Balisacan et al., 2021). While traditional compartmental models provide valuable population-level insights, assumptions of homogeneity limit their ability to represent individual decision-making (Cangiotti et al., 2024; Siegenfeld et al., 2022; Dimarco et al., 2021; Jiang et al., 2024; Bostanci and Conrad, 2025; Rose et al., 2021; Kong et al., 2016). These models also struggle to capture dynamic behavioral changes or adaptive interventions (Brauer et al., 2019; Siegenfeld et al., 2022). Ignoring such factors can limit accuracy and lead to misleading conclusions (Sudhakar et al., 2024; Siegenfeld et al., 2022). Stochastic formulations are particularly relevant for small populations or during the early stages of an outbreak (Brauer et al., 2019).
@@
Agent-Based Models (ABMs). ABMs address the limitations of compartmental models like SIR by modeling heterogeneous individuals using specific attributes and interaction rules within an environment or network (Tracy_Cerdá_Keyes_2018; Hunter et al., 2018; 2017). Furthermore, ABMs excel at capturing individual variability, stochasticity, local interactions, and complex contact networks, which are difficult for compartmental models (Siegenfeld et al., 2022; Hunter et al., 2017). Recent advances demonstrate substantial sophistication such as large-scale ABMs that can simulate millions of agents with GPU acceleration (Chopra et al., 2021) or employ LLM-based adaptive agents (Park et al., 2023).
@@
Challenge of Reward Engineering. Sophisticated ABMs show potential for learning-enabled epidemic modeling. However, systematic evaluation of reward function design remains an important methodological gap. In reinforcement learning, reward design shapes the fundamentals of the learned behavior with different reward structures leading to considerably different learned policies at convergence (Ng et al., 1999; Amodei et al., 2016; Nguyen et al., 2025; Shihab et al., 2025). Despite the importance of this problem in RL, epidemic modeling applications have not systematically investigated how reward structure affects learning under varying conditions.
@@
Reinforcement Learning (RL) for Epidemics. RL has been increasingly leveraged to design policies for epidemic intervention at population and individual level. Individual-based Reinforcement Learning Epidemic Control Agent (IDRLECA) proposed an RL agent at an individual level using GNNs to estimate infection probabilities and balance epidemic suppression vs cost of mobility in its reward function (Feng et al., 2023). VEHICLE (Feng et al., 2022) handles challenges of unobservable asymptomatic cases and delayed effects by using a combination of hierarchical RL with GNNs. Recent work has also focused on using actor critic methods to minimize a cost that considers both the epidemiological, economic and social cost for multi-intervention planning (Mai et al., 2023) or using RL for scheduling lockdown periods (Arango and Pelov, 2020) that incorporates the ICU capacity and minimizing time spent in lockdowns. HRL4EC (Du et al., 2023) also leverages hierarchical RL and multi-component reward that considers both health and socioeconomic cost. SiRL (Bushaj et al., 2023) uses a detailed ABM with RL and feeds the compartmental levels to the agent and uses rewards that are tied to intervention effectiveness and disease spread. These approaches excel at macro-level policy optimization such as determining lockdown time, resource allocation, and intervention sequencing (Mai et al., 2023; Arango and Pelov, 2020; Khadilkar et al., 2020). While valuable for analyzing epidemic policy, these centralized approaches do not examine the impact of different reward formulations on the learned RL policy.
@@ 3 Methodology
ContagionRL is a custom Gymnasium-compatible (Towers et al., 2024) environment that integrates a SIRS+D compartmental epidemiological model (Hethcote, 2000). Built within an agent-based modeling framework, the environment simulates a single reinforcement learning (Sutton and Barto, 2018) agent interacting with a population of humans whose behaviors may be stochastic or deterministic or mix of both. This design enables the study of learned risk-averse navigation and adherence behaviors in response to an evolving epidemic (Saad-Roy and Traulsen, 2023) in different contexts.
Single Agent Design Rationale. Epidemic dynamics are inherently multi-agent interactions (Manfredi and D’Onofrio, 2013; Wang et al., 2016). However, our single-agent approach is a methodological choice that enables transparent attribution of learned behaviors to reward function design, which is the core scientific question of this work. Multi-agent environments introduce complexities such as non-stationarity due to other learning agents (Foerster et al., 2017), emergent social behavior (Leibo et al., 2017) and strategic interactions (Tampuu et al., 2017). These factors obscure whether observed behaviors stem from game-theoretic equilibria or from the reward structure itself. Single-agent formulation permits the identification of essential reward components for effective behavior learning. Understanding individual-level reward-behavior relationships is a necessary prerequisite for designing effective reward systems in multi-agent deployments where each agent still learns from its local reward signal. By isolating a single learning agent in a controlled population of non-learning humans, we can conduct detailed assessments of how different reward formulations affect the learned strategies without interference from adaptive social behaviors. The controlled nature of this methodology allows insights into the relationship between reward and behavior, that will be transferable to more complex multi-agent deployments where understanding individual-level incentive responses becomes critical for designing effective reward systems across a population.
@@ 3.1 SIRS+D Simulation Environment
Environment Dynamics. The simulation happens over discrete timesteps within a two-dimensional grid space of size $G\times G$ with periodic boundary conditions (Allen et al., 2009; Keeling and Rohani, 2008; Frenkel and Smit, 2002). The environment is populated by $N$ non-learning humans and a single controllable Reinforcement Learning (RL) agent. Each individual, human or agent, can exist in one of four states: Susceptible (S), Infected (I), Recovered (R), or Dead (D). We conduct a set of sensitivity analysis in Appendix H and vary the infection rate of disease ($\beta$) in Figure 12 and Table 15, vary the grid size and population density in Figure 13 and Table 16, vary the adherence effectiveness ($\alpha_{\text{eff}}$) in Figure 14 and Table 17, and vary the distance decay of disease transmission in Figure 15 and Table 18.
@@ 3.2 RL Agent Formulation
We implement the agent’s learning problem as a configurable decision-making process that supports both Markov Decision Process (MDP) and Partially Observable MDP (POMDP) formulations. We use both the MDP and POMDP formulation in our experiments. The POMDP employs a threshold-based visibility constraint, where agents can only observe other entities within a specified Euclidean distance radius. The agent’s policy $\pi(a_{t}|o_{t})$, mapping observations to actions, is learned using three deep reinforcement learning algorithms: Proximal Policy Optimization (PPO) (Schulman et al., 2017), Soft Actor-Critic (SAC) (Haarnoja et al., 2018), and Advantage Actor Critic (A2C) a synchronous variant of A3C algorithm (Mnih et al., 2016), all implemented via the stable-baselines3 library (Raffin et al., 2021).
@@
Action Space. At each timestep $t$, the agent selects a continuous-valued action $a_{t}=(\Delta x,\Delta y,\alpha)$, where $(\Delta x,\Delta y)\in[-1,1]^{2}$ defines a movement vector and $\alpha\in[0,1]$ specifies the adherence level to non-pharmaceutical interventions (NPIs). NPIs are methods used to reduce the spread of an epidemic disease without any pharmaceutical drug treatments such as wearing masks or social distancing and are considered extremely effective ways of controlling primary outbreaks that reduce disease transmission and consequentially mortality (Rizvi et al., 2021; Ferguson et al., 2020). The agent’s position is updated using the movement vector, scaled by a maximum step size, and wrapped around the grid’s periodic boundaries:
@@ 3.3 Reward Functions
An important design consideration is the information scope of each reward component. The health and adherence terms depend solely on the agent’s own state and action, making them purely local signals. In contrast, the directional and magnitude components of the potential field reward are computed using the positions and states of all humans in the simulation, regardless of the agent’s visibility radius. This creates a deliberate information asymmetry between the agent’s observations and the reward signal: under POMDP conditions, observations respect the visibility constraint (Section 3.2), while the reward provides a globally informed training signal. This design functions as a form of privileged critic (Pinto et al., 2017), where the reward shapes policy gradients using complete state information during training, but the learned policy must generalize from local observations alone at evaluation time. Table 6 summarizes the information scope of each reward component.
@@ 4 Baselines
Greedy Distance Maximizer Baseline. This heuristic policy implements a specific survival strategy focused on immediate risk avoidance. At each step, it first identifies the nearest infected human. It then evaluates a predefined set of potential movement actions across the eight cardinal and intercardinal directions, plus staying stationary. It systematically simulates the outcome of each potential action, by computing the hypothetical distance to the nearest infected human if that move were taken, and selects the action that yields the greatest distance separation. The policy’s evaluation of counterfactual outcomes provides a situational awareness absent in simpler baselines. While myopic and handcrafted, this policy provides a strong heuristic benchmark by explicitly evaluating the spatial consequence of each potential action.